OpenStudy (anonymous):

What is the percent yield of ferrous sulfide if the actual yield is 220.0 g and the theoretical yield is 275.6 g? 12.5% 55.6% 79.8% 87.9%

OpenStudy (anonymous):

\[\color{green}{\begin{align*}\% \text{yield} &= \frac{\text{actual yield}}{\text{theoretical yield}} \times 100 \\ \\&= \frac{222.0~\text{g}}{275.6~\text{g}} \times 100 \\ &\text{etc.} \end{align*}}\]